2. Simplify the following expression and provide the truth table (i) A. B + A.B (ii) (Ā+Ā)
Answer #1

kindly see the attached image for the solution

(1) A.Bt A.B = A (B+B) = A (1) (since Bt B 21) Truth table ABB A.BLA.BA. B+ A.B - oo Same as A. (c) (A+) = (a) (sme pa tā 27)